A wire oriented in the east-west direction carries a current east -ward.Direction of the magnetic field at a point to the south of wire wil be?

Grade:12th pass

1 Answers

Susmita
5 years ago
Allright.So draw x-y axis on your paper.The x axis is the wire extending from negative x axis to positive x axis.The current is flowing from negative x axis to positive x axis.We are asked to find out direction of B at any point on negative y axis(or parallel to x axis in south).
Use the right hand screw rule.You will see B is into the plane of the paper,ie,along the negative z axis.
If you want to find it mathematically,use the biosavart law
B=\frac{\mu}{4\pi}\int \frac{\vec{I}dl*\vec{l}}{l^3}
Here Vec I=vec x and vec l=vec (-y)
The cross product Idl×l will have a direction 
vec x×vec(-y)=-vec z